OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 18mthe square root of 5 times m. and width the square root of 4 times m., which of the following is the area of the rectangle?

OpenStudy (anonymous):

\[(18\sqrt5\times m)\times(\sqrt4 \times m)\] Is this the question? (As in, do the length and width look right to you?)

OpenStudy (anonymous):

\[\frac{ 18m \sqrt{5m} }{ \sqrt{4m} }\]

OpenStudy (anonymous):

oops you have to multipy them to find the area..

OpenStudy (anonymous):

like the top and bottom

OpenStudy (anonymous):

\[18m \sqrt{5m} \times \sqrt{4m}\]

OpenStudy (anonymous):

is m the meter or is m an unknown??

OpenStudy (anonymous):

its just a variable i guess

OpenStudy (anonymous):

\[18m \sqrt{5m} \times 2 \sqrt m\]

OpenStudy (anonymous):

can you solve from here?

OpenStudy (anonymous):

\[36m^2 \sqrt5\]

OpenStudy (anonymous):

\[36m \sqrt{5m^2}\] (The step before)

OpenStudy (anonymous):

thank you so much :))))

OpenStudy (anonymous):

you're welcome.